OpenStudy (anonymous):

The difference of x and a number is 6. What is the other number?

OpenStudy (jdoe0001):

http://www.mathsisfun.com/numbers/subtraction.html

OpenStudy (jdoe0001):

ohh,, nevermind, misread well just solve for "x" x- n = 6 # substract "x" to both sides x - n -x = 6 -x -n = 6 - x #multiply both sides by -1 -1(-n) = -1(6-x) n = x+6